OpenStudy (anonymous):

please help me... how to answer this question in substitution? 2x+3y=7 5x-2y=8

OpenStudy (anonymous):

Choose one of the equations and calculate y. Sub the value of y in the other equation to find x. When u have x sub it back in an equation to get y.

OpenStudy (anonymous):

We'll determine x from the top equation. 2x = 7 - 3y x = (7-3y)/2 Now, we'll replace x from the 2nd equation by the expression (7-3y)/2 5*(7-3y)/2 - 2y = 8 We'll remove the brackets: 35 - 15y - 4y = 16 -19y = 16 - 35 -19y = -19 y = 1 x = (7-3)/2 x = 2

OpenStudy (anonymous):

thank you very much 4 the answer my fren...